OpenStudy (anonymous):

A student is assessing the correlation between the average number of hours of internet browsing and average score on a math test for the students of a class. The table below shows the data: Number of hours of internet browsing (x) 0 0.5 1 1.5 2 2.5 3 3.5 4 4.5 Score on math test (y) 95 94 93 92 91 90 89 87 86 85 Part B: Write a function which best fits the data.

OpenStudy (anonymous):

@Qwertty123

OpenStudy (qwertty123):

Do you have a like TI83 or TI84 calculator?

OpenStudy (anonymous):

no :/ dont even know what that is

OpenStudy (qwertty123):

oh okay.. http://www.alcula.com/calculators/statistics/linear-regression/

OpenStudy (anonymous):

i have a guess dunno if its right. 1/0.5 = 2 so that is slope. y-1=2(x-0.5)

OpenStudy (anonymous):

@Qwertty123 ^